# RAG-Enhanced Collaborative LLM Agents for Drug Discovery

### v2

@@ 3.2.2 Knowledge Graph Team
Importantly, while both the DrugRel Agent and BioRel Agent aim to reason about the query molecule in relation to other relevant drugs in the KG for the specific task, they leverage distinct knowledge sources and perform different roles. Specifically, the BioRel Agent primarily leverages the connectivity between drugs and other biological entities in the KG, focusing on interpreting and summarizing this network of relationships and aligning it with the specific task at hand. In contrast, the DrugRel Agent primarily draws on its internal knowledge, triggered by the names of the related drug entities in the KG, and incorporates structural similarity between them. In Section 4, we demonstrate how these agents complement each other effectively, producing a synergistic effect when combined together.
3.2.3 Molecular Understanding Team
While the KG Team compiles the report by aggregating contextual knowledge, the Molecule Understanding (MU) Team focuses primarily on the query molecule itself. The MU Team is composed of a single Molecule Understanding (MU) Agent, which aims to write a report on the query molecule by leveraging its structural information, annotations from tools, and reports from other agents.
Molecule Understanding (MU) Agent. The MU Agent retrieves annotations for the query molecule, denoted as $c_{q}$. If the Planning Team decides to use external annotation tools (i.e., $o_{\text{MAP}}=\text{TRUE}$), additional captions $\tilde{c}_{q}$ are generated with the external captioning tools as follows:
$\small\tilde{c}_{q}=\text{Captioning Tools}(g_{q}),\vspace{-5pt}$ (6)
and concatenated to the annotations retrieved from the database: $c_{q}=c_{q}||\tilde{c}_{q}$. External captioning tools allow the system to easily harness recent advances in LLM-driven molecular understanding (Pei et al., 2023; Yu et al., 2024), and can potentially include any specialized tool.
The agent then analyzes the structure of the molecule, contextualizing it with reports generated by the other agents. Using the SMILES representation, the caption of the query molecule, and the reports from the KG Team, it compiles a comprehensive molecular annotation report as follows:
$\small o_{\text{MUA}}=\text{MU Agent}(g_{q},c_{q},o_{\text{DRA}},o_{\text{BRA}},\mathcal{I}).\vspace{-5pt}$ (7)
3.2.4 Prediction Agent
@@
$\small\mathcal{A}_{g_{q}}=\text{Task Agent}(g_{q},o_{\text{MUA}},o_{\text{DRA}},o_{\text{BRA}},\mathcal{I}).\vspace{-5pt}$ (8)
By integrating this evidence, the Prediction Agent can perform a comprehensive analysis of the query molecule. Importantly, the output of the Prediction Agent can be flexibly adjusted based on the specific task requirements. For instance, it can be a descriptive caption, a simple yes/no response for binary classification, or a more complex answer listing the top-k targets associated with the query molecule. Such behavior leverages the zero-shot capabilities of LLMs (Kojima et al., 2022) and does not require additional fine-tuning. Therefore, a key advantage of CLADD is its flexibility, which enhances scientist-AI interactions.
@@ 4 Experiments
We assess the effectiveness of CLADD by conducting a range of drug discovery tasks, including property-specific molecular captioning (Section 4.1), drug-target prediction (Section 4.2), and drug toxicity prediction (Section 4.3).
Implementation Details. In all experiments, we utilize GPT-4o mini through the OpenAI API for each agent. We use PrimeKG (Chandak et al., 2023) as the KG, PubChem (Kim et al., 2021) as an annotation database, and MolT5 (Edwards et al., 2022) as an external captioning tool. Additional implementation details and agent templates can be found in Appendix E and G, respectively.
4.1 Property-Specific Molecular Captioning Task
Earlier studies on molecular captioning tasks have primarily focused on generating general descriptions of molecules without targeting specific areas of interest, raising concerns about their practical applicability in real-world drug discovery tasks. Indeed, the usefulness of a molecular description is often task-dependent, and scientists may be interested in detailed explanations of specific characteristics of a molecule rather than a general description (Guo et al., 2024; Edwards et al., 2024). Hence, in this paper, we introduce property-specific molecular captioning, where the model is required to generate a description for a given molecule customized to a particular area of interest.
Datasets. We leverage four widely recognized molecular property prediction datasets from the MoleculeNet benchmark (Wu et al., 2018): BBBP, Sider, ClinTox, and BACE. We provide further details on the datasets in Appendix C.
Methods Compared. We consider different baseline approaches. First, we compare recent molecular captioning methods designed to generate general descriptions of molecules, including MolT5 (Edwards et al., 2022), LlasMol (Yu et al., 2024), and BioT5 (Pei et al., 2023). Furthermore, we assess general-purpose LLMs, namely GPT-4o mini and GPT-4o. Finally, we consider two GNNs pre-trained with different methodologies: GraphMVP (Liu et al., 2021) and MoleculeSTM (Liu et al., 2023). We provide further details on the baseline models in Appendix D.
BBBP Sider ClinTox BACE
GNNs
GraphMVP 69.59 (1.29) 60.88 (0.41) 87.57 (3.26) 80.24 (2.92)
MoleculeSTM 70.14 (0.90) 58.69 (0.89) 92.19 (2.79) 79.24 (3.40)
Only SMILES 70.95 (1.14) 60.80 (1.18) 91.62 (2.18) 74.21 (1.32)
General LLMs
GPT-4o mini 67.85 (1.50) 58.18 (1.55) 90.74 (1.91) 74.22 (1.95)
GPT-4o 66.43 (1.47) 60.41 (1.21) 88.13 (1.74) 67.82 (4.14)
Domain LMs
MolT5 69.77 (1.89) 57.20 (0.98) 87.91 (1.25) 74.28 (4.00)
LlasMol 68.12 (1.48) 61.50 (1.66) 89.67 (0.57) 75.42 (2.98)
BioT5 69.68 (1.23) 64.65 (2.01) 92.80 (2.92) 77.23 (1.95)
CLADD 72.28 (1.04) 66.42 (1.31) 93.80 (2.30) 77.74 (3.15)
Evaluation Protocol. Although property-specific captions are practical, no ground truth property-specific captions exist for individual molecules, rendering traditional text generation evaluation methods inapplicable. Thus, in line with recent works (Xu et al., 2024; Guo et al., 2024; Edwards et al., 2024), we assess whether the generated captions can drive a classification model that categorizes molecules based on their properties. Specifically, for a generated caption and the SMILES representation of the target molecule, we concatenate them using a [CLS] token, forming SMILES[CLS]caption, and fine-tune a SciBERT (Beltagy et al., 2019) model for property prediction. The “Only SMILES” model utilizes only the SMILES string as input for the SciBERT classifier. For baseline GNNs, we convert the SMILES into a molecular graph and provide it to the model. For all the experiments, we use a scaffold splitting strategy to simulate realistic distribution shifts (train/validation/test data split as 80/10/10%), following previous works (Liu et al., 2023). We perform five independent fine-tuning runs of SciBERT (or GNN baselines) and report the mean and standard deviation of the AUROC.
Experimental Results. Table 1 summarizes the results. We note the following observations: 1) While domain-specific models outperform general-purpose LLMs, their performance remains suboptimal, occasionally falling behind the “Only SMILES” approach. This means that the generated captions occasionally reduce model performance compared to using only the SMILES representation of the molecule. This aligns with previous work that found that general descriptors may lack property-specific relevance (Guo et al., 2024; Edwards et al., 2024). 2) On the other hand, CLADD-generated captions consistently outperform all the baseline captioners and successfully improve over “Only SMILES” across all datasets. We attribute this improvement to the ability of CLADD to draw on external biochemical knowledge to ground its generation and its task-specificity. 3) Moreover, CLADD consistently outperforms pre-trained GNN baselines, except on the BACE dataset. Interestingly, this is also the only dataset for which the “Only SMILES” baseline falls short compared to GNN models, thus highlighting the critical role of 2D topological and 3D geometric information in this case. This paves the way for future research on injecting essential aspects of molecules, such as topological and geometric information, into LLM understanding.
Table 2: Performance in drug target tasks (Precision @ 5). Bold and underline indicate best and second-best language model-based methods.
@@ 4.2 Drug-Target Prediction Task
Accurately predicting a drug’s protein target is essential for understanding its mechanism of action and optimizing its therapeutic efficacy while minimizing off-target effects (Santos et al., 2017; Batool et al., 2019). Here, we evaluate the models’ ability to accurately identify which proteins a given molecule is most likely to activate or inhibit.
Datasets. We use molecular targets present in the Drug Repurposing Hub (Corsello et al., 2017), DrugBank (Wishart et al., 2018), and STITCH v5.0 (Szklarczyk et al., 2016), as preprocessed in Zheng et al., 2023, including 13,688 molecules in total (details are presented in Appendix C).
Methods Compared. We evaluate two pre-trained GNNs, GraphMVP and MoleculeSTM, along with two general-purpose LLMs—GPT-4o mini and GPT-4o, and the domain-specific language model Galactica (Taylor et al., 2022) (details are presented in Appendix D).
Evaluation Protocol. We assess the performance of LLMs in a zero-shot setting. Specifically, for a given target molecule, we prompt the LLMs to generate the top 5 proteins that the molecule is most likely to activate or inhibit. The precision is then calculated by determining whether the generated proteins match the correct answers. As baseline GNNs cannot perform this task without training in a zero-shot setting, we fine-tune them in a few-shot setting using 10% of the data. For domain-specific LMs, we also present fine-tuning results on the specific task. To better assess generalization power, we separately report the performance on the test set for molecules present/not present in the external databases (“Overlap”/“No Overlap”).
Table 3: Performance in drug toxicity prediction task (Macro-F1). Avg. indicates the average performance over four datasets. Bold and underline indicate best and second-best methods.
hERG DILI Skin Carcinogens Avg.
@@
GPT-4o mini 28.42 33.47 41.84 69.51 43.31
GPT-4o 40.45 25.76 54.51 38.91 39.90
Domain LMs
Galactica 125M 40.78 33.56 42.43 17.64 33.60
Galactica 1.3B 48.57 34.37 42.43 44.00 42.34
Galactica 6.7B 23.75 57.67 40.41 65.20 46.76
GIMLET 36.50 35.51 42.28 48.47 40.69
LlasMol 23.75 61.20 31.92 51.68 42.14
CLADD 51.46 41.10 50.43 58.31 50.33
Experimental Results. Table 2 summarizes the results. We observe the following: 1) CLADD outperforms all the baselines, with a higher likelihood of correctly identifying proteins activated/inhibited by the input molecule. 2) Importantly, the superiority of CLADD is confirmed for molecules not present in the caption database or knowledge graph (Table 2 (b)), showcasing CLADD’s ability to leverage external knowledge to generalize to novel molecules. This outcome underscores the potential of CLADD for exploring entirely new molecular spaces. 3) We observe that domain-specific fine-tuned models, such as Galactica, GIMLET, and MolecularGPT, could not generate five protein targets in a zero-shot setting when prompted to do so, as they are limited to providing answers based on their fine-tuning instruction dataset. By specifically fine-tuning Galactica on the task, we were able to answer the specific question, outperforming general-purpose LLMs in most experiments, but results were still inferior to CLADD. This further highlights the flexibility of CLADD, which leverages the zero-shot abilities of general-purpose LLMs in its architecture.
@@
Figure 3: Example of collaboration between agents in CLADD (on the drug-target prediction task). Red represents adrenergic receptors, yellow represents histamine receptors, and green represents dopamine receptors. The full version is available in Appendix F.
4.3 Drug Toxicity Prediction Task
Accurate predictions of drug toxicity are crucial to ensure patient safety and minimize the risk of adverse effects during drug development (Basile et al., 2019). Here, we evaluate the models’ ability to predict the toxicity of a target molecule from its SMILES-based structural description.
Datasets. We use four datasets to comprehensively evaluate the performance of CLADD for drug toxicity prediction tasks: hERG (Wang et al., 2016), DILI (Xu et al., 2015), Skin (Alves et al., 2015), Carcinogens (Lagunin et al., 2009) datasets (details are presented in Appendix C).
Methods Compared. We compare five domain-specific LLMs—Galactica 125M, Galactica 1.3B, Galactica 6.7B (Taylor et al., 2022), LlasMol (Yu et al., 2024), and GIMLET (Zhao et al., 2023a) alongside two general-purpose LLMs, GPT-4o and GPT-4o mini (details in Appendix D).
Evaluation Protocol. As all the methods compared are foundation models, we evaluate their performance in a zero-shot setting. Specifically, given a SMILES-based structural description of the target molecule and a task description, the model outputs whether the molecule possesses the target property (binary classification). Using the text-formatted output generated by each model, we compute the Macro-F1 score (Opitz & Burst, 2019) as the evaluation metric.
Experimental Results. Table 3 summarizes the results. 1) Comparing Galactica models, we observe that performance improves as the model size increases. However, this trend is not confirmed for GPT-4o and GPT-4o mini. This suggests that larger general-purpose models do not necessarily excel in domain-specific tasks, a trend also reported in prior research (Edwards et al., 2024). These findings emphasize the necessity of domain-specific training in combination with model scaling. 2) Meanwhile, CLADD outperforms all the baselines (average score across datasets), without requiring domain-specific training by effectively incorporating external knowledge into general-purpose LLMs. This highlights the potential of integrating external knowledge at inference time to improve molecular understanding, acting as an alternative (or complementary approach) to expensive domain-specific model fine-tuning.

### v3

@@ 3.1 CLADD
Importantly, while both the DrugRel Agent and BioRel Agent aim to reason about the query molecule in relation to other relevant entities in the KG for the specific task, they leverage distinct knowledge sources and perform different roles. Specifically, the BioRel Agent focuses on summarizing the network of relationships between drugs and other biological entities in the KG, contextualizing it with respect to the specific task at hand. In contrast, the DrugRel Agent primarily draws on its internal knowledge, triggered by the names of the related drug entities in the KG, and incorporates structural similarity between them. In Section 4, we demonstrate how these agents complement each other, producing a synergistic effect when combined together.
Molecular Understanding Team
The Molecular Understanding (MU) Team compiles a report on the query molecule by leveraging external annotations and integrating them with structural information and reports from other agents.
Molecule Annotations. Annotations from the external database are retrieved for the query molecule, denoted as $c_{q}$. If the Planning Team decided to use external annotation tools (i.e., $o_{\text{MAP}}=\text{TRUE}$), additional captions $\tilde{c}_{q}$ are generated with the external captioning tools as follows:
$\tilde{c}_{q}=\text{Captioning Tools}(g_{q}),$ (5)
and concatenated to the annotations retrieved from the database: $c_{q}=c_{q}||\tilde{c}_{q}$. External captioning tools allow the system to easily harness recent advances in LLM-driven molecular understanding (pei2023biot5; yu2024llasmol), and can potentially include any tools, given that the output can be transformed into text.
Molecule Understanding (MU) Agent. The MU agent then analyzes the structure of the molecule, combining it with annotations and reports generated by the KG Team and generating a comprehensive report as follows:
$o_{\text{MUA}}=\text{MU Agent}(g_{q},c_{q},o_{\text{DRA}},o_{\text{BRA}},\mathcal{I}).$ (6)
Prediction Agent
@@
$\mathcal{A}_{g_{q}}=\text{Task Agent}(g_{q},o_{\text{MUA}},o_{\text{DRA}},o_{\text{BRA}},\mathcal{I}).$ (7)
By integrating this evidence, the Prediction Agent can perform a comprehensive analysis of the query molecule. Importantly, the output of the Prediction Agent can be flexibly adjusted based on the specific task requirements. For instance, it can be a descriptive caption, a simple yes/no response for binary classification, or an open-ended answer. Such behavior leverages the zero-shot capabilities of LLMs (kojima2022large) and does not require additional fine-tuning. Therefore, a key advantage of CLADD is its flexibility, which enhances scientist-AI interactions.
@@ 4 Experiments
Implementation Details. In all experiments, we utilize GPT-4o mini through the OpenAI API for each agent. We use PrimeKG (chandak2023building) as the KG, PubChem (kim2021pubchem) as an annotation database, and MolT5 (edwards2022translation) as an external captioning tool. Additional implementation details and agent templates can be found in Appendix E and G, respectively.
@@
Table 1: Performance in drug-target prediction tasks (Precision @ 5). Bold and underline indicate best and second-best language model-based methods.
@@ 4.1 Drug-Target Prediction Task
Accurately predicting a drug’s protein target is essential for understanding its mechanism of action and optimizing its therapeutic efficacy while minimizing off-target effects (santos2017comprehensive; batool2019structure). Here, we evaluate the models’ ability to accurately identify which proteins a given molecule is most likely to activate or inhibit in a set prediction setting.
Datasets. We use molecular targets present in the Drug Repurposing Hub (corsello2017drug), DrugBank (wishart2018drugbank), and STITCH v5.0 (szklarczyk2016stitch), as preprocessed in zheng2023chempert, including 13,688 molecules in total (details are presented in Appendix C).
Methods Compared. We evaluate two pre-trained GNNs, GraphMVP and MoleculeSTM, along with two general-purpose LLMs—GPT-4o mini and GPT-4o, and the domain-specific language model Galactica (taylor2022galactica) (details are presented in Appendix D).
Evaluation Protocol. We assess the performance of LLMs in a zero-shot setting. Specifically, for a given target molecule, we prompt the LLMs to generate the top 5 proteins that the molecule is most likely to activate or inhibit, and we calculate the precision with respect to ground truth data. As baseline GNNs cannot perform this task without training in a zero-shot setting, we fine-tune them in a few-shot setting using 10% of the data. For domain-specific LMs, we also present fine-tuning results on the specific task. To better assess generalization power, we separately report the performance on the test set for molecules present/not present in the external databases (“Overlap”/“No Overlap”).
Experimental Results. Table 1 summarizes the results. We observe the following: 1) CLADD outperforms all the baselines, with a higher likelihood of correctly identifying proteins activated/inhibited by the input molecule. 2) Importantly, the superiority of CLADD is confirmed for molecules not present in the caption database or knowledge graph (Table 1 (b)), showcasing CLADD’s ability to leverage external knowledge to generalize to novel molecules. 3) We observe that domain-specific fine-tuned models, such as Galactica, GIMLET, and MolecularGPT, could not perform this task in a zero-shot setting when prompted to do so, likely because this task is not included in their fine-tuning instruction dataset. By specifically fine-tuning Galactica on the task, we were able to answer the specific question, outperforming general-purpose LLMs in most experiments, but results were still inferior to CLADD. This further highlights the flexibility of CLADD, which leverages the zero-shot abilities of general-purpose LLMs in its architecture.
BBBP Sider ClinTox BACE
GNNs
GraphMVP 69.59 (1.29) 60.88 (0.41) 87.57 (3.26) 80.24 (2.92)
MoleculeSTM 70.14 (0.90) 58.69 (0.89) 92.19 (2.79) 79.24 (3.40)
Only SMILES 70.95 (1.14) 60.80 (1.18) 91.62 (2.18) 74.21 (1.32)
@@
GPT-4o mini 67.85 (1.50) 58.18 (1.55) 90.74 (1.91) 74.22 (1.95)
GPT-4o 66.43 (1.47) 60.41 (1.21) 88.13 (1.74) 67.82 (4.14)
Domain LLMs
MolT5 69.77 (1.89) 57.20 (0.98) 87.91 (1.25) 74.28 (4.00)
LlasMol 68.12 (1.48) 61.50 (1.66) 89.67 (0.57) 75.42 (2.98)
BioT5 69.68 (1.23) 64.65 (2.01) 92.80 (2.92) 77.23 (1.95)
CLADD 72.28 (1.04) 66.42 (1.31) 93.80 (2.30) 77.74 (3.15)
4.2 Property-Specific Molecular Captioning Task
Earlier studies on molecular captioning tasks have primarily focused on generating general descriptions of molecules without targeting specific areas of interest, raising concerns about their practical applicability in real-world drug discovery tasks. Indeed, the usefulness of a molecular description is often task-dependent, and scientists may be interested in detailed explanations of specific characteristics of a molecule rather than a general description (guo2024moltailor; edwards2024molcap). Hence, in this paper, we introduce property-specific molecular captioning, where the model is required to generate a description for a given molecule customized to a particular task of interest.
Datasets. We leverage four widely recognized molecular property prediction datasets from the MoleculeNet benchmark (wu2018moleculenet): BBBP, Sider, ClinTox, and BACE (further details in Appendix C).
Methods Compared. We consider different baseline approaches. First, we compare recent molecular captioning methods designed to generate general descriptions of molecules, including MolT5 (edwards2022translation), LlasMol (yu2024llasmol), and BioT5 (pei2023biot5). Furthermore, we assess general-purpose LLMs, namely GPT-4o mini and GPT-4o. Finally, for reference, we consider standard molecular property prediction baselines, including two GNNs pre-trained with different methodologies: GraphMVP (liu2021pre) and MoleculeSTM (liu2023multi). We provide further details on the baseline models in Appendix D.
(a) Toxicity (b) MLSMR
hERG DILI Skin Avg. Mtb
General LLMs
GPT-4o mini 28.42 33.47 41.84 34.58 33.33*
GPT-4o 40.45 25.76 54.51 40.24 36.68
Domain LLMs
Galactica 125M 40.78* 33.56 42.43 38.92 33.33*
Galactica 1.3B 48.57 34.37 42.43 41.79 33.33*
Galactica 6.7B 23.75* 57.67 40.41* 40.61 33.33*
GIMLET 36.50 35.51 42.28 38.09 39.81
LlasMol 23.75* 61.20 31.92 38.95 33.33*
CLADD 51.46 41.10 50.43 47.66 50.92
Table 3: Performance in biological activity prediction task including (a) toxicity and (b) antibacterial activity (Macro-F1). Avg. indicates the average performance over toxicity datasets. Bold and underline indicate best and second-best methods. * indicates whether the model always outputs the same response, either “Yes” or “No”.
Evaluation Protocol. Although property-specific captions are practical, no ground truth property-specific captions exist for individual molecules, rendering traditional text generation evaluation methods inapplicable. Thus, in line with recent works (xu2024llm; guo2024moltailor; edwards2024molcap), we assess whether the generated captions can drive a classification model that categorizes molecules based on their properties. Specifically, we pose this evaluation as a molecular property prediction problem, and fine-tune a SciBERT model (beltagy2019scibert) on the generated caption concatenated to the SMILES representation to predict the property of interest. The “Only SMILES” model utilizes only the SMILES string as input for the SciBERT classifier. For baseline GNNs, each SMILES string is converted into a molecular graph. For all the experiments, we use a scaffold splitting strategy to simulate realistic distribution shifts, following previous work (liu2023multi). This evaluation protocol is further illustrated in Appendix C.2.
Experimental Results. Table 2 summarizes the results. 1) While domain-specific LLMs outperform general-purpose LLMs, their performance remains suboptimal, occasionally falling behind the “Only SMILES” approach. This means that the generated captions occasionally reduce model performance compared to using only the SMILES representation of the molecule. This aligns with previous work that found that general descriptors may lack property-specific relevance (edwards2024molcap). 2) On the other hand, CLADD-generated captions consistently outperform all the baseline captioners and successfully improve over “Only SMILES” across all datasets. We attribute this improvement to the ability of CLADD to draw on external biochemical knowledge to ground its generation and its task-specificity. 3) Moreover, CLADD consistently outperforms pre-trained GNN baselines, except on the BACE dataset. Interestingly, this is also the only dataset for which the “Only SMILES” baseline falls short compared to GNN models, thus highlighting the critical role of 2D topological and 3D geometric information in this case. This paves the way for future research on injecting essential aspects of molecules, such as geometric information, into LLM understanding.
@@ 4.2 Property-Specific Molecular Captioning Task
Figure 3: Example of collaboration between agents in CLADD (on the drug-target prediction task). Red represents adrenergic receptors, yellow represents histamine receptors, and green represents dopamine receptors (The full version in Appendix F).
4.3 Biological Activity Prediction
Accurately predicting molecular bioactivity is a cornerstone of drug discovery, which is often hindered by the existence of countless biological contexts and sparse experimental data. We therefore explore the zero-shot characterization of biological activity for unseen compounds. To this goal, we focus on drug toxicity (basile2019artificial) and antibacterial activity (melo2021accelerating) prediction.
Datasets. For drug toxicity prediction, we use three benchmark datasets: hERG (wang2016admet), DILI (xu2015deep), and Skin (alves2015predicting). For antibacterial activity prediction, we use a dataset curated from eke2025genetic, hereafter referred to as MLSMR_Mtb. In addition to its relevance, we selected MLSMR_Mtb for its recency, as it was published after GPT-4o training and in parallel to the preparation of this study, therefore avoiding the risk of pre-training data leakage. Dataset details are presented in Appendix C.
Methods Compared. We compare five domain-specific LLMs—Galactica 125M, Galactica 1.3B, Galactica 6.7B (taylor2022galactica), LlasMol (yu2024llasmol), and GIMLET (zhao2023gimlet), alongside two general-purpose LLMs, GPT-4o and GPT-4o mini (details in Appendix D).
Evaluation Protocol. Evaluation follows a zero-shot QA setting. The input includes a SMILES representation of the molecule and the task description. Using the text-formatted output generated by each model, we compute the Macro-F1 score (opitz2019macro) as the evaluation metric.
Experimental Results. Table 3 summarizes the results. 1) Both on toxicity datasets (average score) and the recently published antibacterial activity dataset, CLADD outperforms all the baselines. This includes GPT-4o mini, which is used as building block of CLADD. This highlights its ability to perform zero-shot predictions without domain-specific fine-tuning by effectively incorporating external knowledge into general-purpose LLMs at inference time. 2) Notably, for three datasets (hERG, Skin and MLSMR_Mtb), several baseline models often output the same response, either “Yes” or “No”, indicating their inability to perform the given task. In contrast, CLADD did not suffer from this limitation. 3) Most baselines performed especially poorly on the recently released dataset (MLSMR_Mtb). However, CLADD shows a significant improvement over the baselines, demonstrating its reasoning ability on unseen tasks that are even published after the language model.